OpenStudy (anonymous):

In the spring, the height of a vine was 25 inches. In the fall, the height of the vine was 140 inches. What was the percent of increase? 46% 56% 460% 560% I need to ace this test and get my grade up! Thanks for the help if u can!

OpenStudy (anonymous):

Would the answer be 46%?

OpenStudy (phi):

what do you get for change= end height - start height?

OpenStudy (anonymous):

115 = 140 - 25

OpenStudy (phi):

and what fraction do you get doing change/start ?

OpenStudy (anonymous):

115/25 = 4.6 ?

OpenStudy (phi):

yes, an "improper fraction" because it's bigger than 1. but that's ok. finally, how do you change fraction to a percent ? multiply by 100 and add a % sign to the end

OpenStudy (anonymous):

0.046 ???

OpenStudy (phi):

you divided. you want to multiply. example: 1/2 = 0.5 multiply by 100 to get 50 (and add %) to get 50% 1/2 is 50%

OpenStudy (anonymous):

no 460% ?

OpenStudy (anonymous):

so 460% right?

OpenStudy (phi):

if you get confused about divide/multiply think: fraction = x/100 so (multiply both sides by 100) 100*fraction = x if that makes sense?? but yes, 460%

OpenStudy (triciaal):

yes

OpenStudy (anonymous):

Thank you! Okay here is a second question. What is 0.1% written as a decimal?

OpenStudy (phi):

or if you get confused, do a simple problem. You know 1/2 = 50% = 50 /100 also 1/2 = 0.5 so it's common sense that to get from 0.5 to 50, you multiply by 100

OpenStudy (anonymous):

Ok awesome thanks

OpenStudy (phi):

x% means (by definition) x/100

OpenStudy (phi):

so rewrite 0.1% as 0.1 divided by ?

OpenStudy (anonymous):

wait i think i figured it out. would the answer be 0.001?

OpenStudy (phi):

yes. you could say 0.1% is 0.1/100 you can use a calculator, or you can say .1 is 1/10 so .1/100 is 1/10 * 1/100 = 1/1000 and hopefully you can write that as 0.001

OpenStudy (anonymous):

twenty is what percent of 400? my answer would be 80%. this is correct right?

OpenStudy (phi):

20 out of 400? write it as a fraction

OpenStudy (anonymous):

wait no! it would be 5% hahahhaa

OpenStudy (phi):

yes. It helps to think about "real world" somebody owes you $400. they give you $20 and say they paid you 80%. would you believe them ?

OpenStudy (anonymous):

right. thanks. Heres the last one! Ruiz invests $550 at a simple interest rate of 4%. How much interest will he earn after 5 years?

OpenStudy (phi):

any idea what's 4% of 550 ?

OpenStudy (phi):

change 4% to a fraction, and multiply that fraction times 550

OpenStudy (phi):

4% means 4 per cent which is Latin for 4 per 100

OpenStudy (phi):

or 4 out of 100

OpenStudy (phi):

if you can figure out 4% of 550, that is the interest you get for 1 year multiply by 5 to get how much interest for 5 years.

OpenStudy (anonymous):

okay. would the answer be $110 ?

OpenStudy (phi):

what's 4% of 550 ?

OpenStudy (anonymous):

22?

OpenStudy (phi):

yes. that is the interest for 1 year (4% per year) now multiply by 5 to get the interest for 5 years

OpenStudy (anonymous):

and so 22 x 5 is 110

OpenStudy (phi):

yes. It looks like 110 is the answer.

OpenStudy (phi):

good job

TheSmartOne (thesmartone):

Formula for simple interest: \(\Large I = Prt\) \(\large Where:\\\ I =~ interest\\\ P = ~amount ~borrowed~ (called~ "Principal")\\\ r =~ interest~ rate\\\ t =~time \) just in case you wanted to know the formula for it :)

OpenStudy (anonymous):

Thanks!!

OpenStudy (anonymous):

i aced my test. thanks guys!

TheSmartOne (thesmartone):

also note that r has to be in decimal and t is in years.

TheSmartOne (thesmartone):

O_O test?

TheSmartOne (thesmartone):

Ummm. Did you know that OS has an anti-cheating policy?

TheSmartOne (thesmartone):

And that your school has an honor code?

TheSmartOne (thesmartone):

Anti-Cheating Policy What is OpenStudy's Anti-Cheating Policy? Openstudy.com is proud to be a pioneer in the online education community and is dedicated to advancing knowledge, understanding and participation in student success. Through OpenStudy.com, students, educators and subject enthusiasts from around the globe participate in a dynamic social learning experience that stands above the rest. The resources we provide, study groups and Users, are here to help You study smarter and study with friends. We believe You will actually enjoy learning now! We work diligently to prevent misuse, but ultimately, You make the choice to use or misuse the resources we provide. By using the OpenStudy Site or Service, you agree to abide by OpenStudy’s Anti-cheating Policy and to also comply with copyright laws and the academic policies or honor codes of your school and course, and to report violations promptly to us. Posting questions from tests, exams or other third party content may be an infringement of copyrights of the owner of the posted content. In most cases, asking for help on an exam or test, or helping someone during their test or exam, will be considered cheating by the institution or by OpenStudy. OpenStudy reserves the right to remove such content and/or suspend, ban or otherwise remove Users from the system who are suspected of cheating. If you feel there is an error, You may email contact@openstudy with an explanation but You acknowledge that the determination to remove content or suspend, ban or otherwise remove a User from the system is within the sole discretion of Company. By using the OpenStudy Site or Service, you acknowledge that OpenStudy cooperates fully with schools, colleges and other educational institutions (together “Schools”) and reserves the right to disclose User information to a School if there is an allegation by that School of cheating or violation of the School’s academic policies. Openstudy.com will take swift action against anyone found misusing the community’s Site or Service. In addition to refunding any membership fee, we will ban Users who cheat from the Site and Service permanently. A message to students:Please also keep in mind that Openstudy is a community for educators as well as students. There are many educators, school and college administrators and parents on the system. Your questions and answers on OpenStudy are archived and openly visible to everyone and served up by Google and other searchbots. As in any social network, your digital history is open and visible for years to come. If you cheat when using OpenStudy, it will be visible, and retrievable evidence. We want you to make the responsible and ethical choice. We require every member to abide by OpenStudy’s Anti-cheating Policy. Openstudy.com will take swift action against anyone found misusing the community’s site. In addition to refunding any membership fee, we will ban users who cheat from the site permanently. If you believe there has been a violation of academic integrity on our site, please do not hesitate to contact us at contact@openstudy.com I agree to abide by the AntiCheating policies. From: http://openstudy.com/terms-and-conditions

OpenStudy (anonymous):

lol pre test ! it just has similar questions to the actual test

TheSmartOne (thesmartone):

oh lol

OpenStudy (anonymous):

haha yeah i wouldn't cheat

TheSmartOne (thesmartone):

Well anyhoot @bean73 \(\bf\huge\color{#ff0000}{W}\color{#ff2000}{e}\color{#ff4000}{l}\color{#ff5f00}{c}\color{#ff7f00}{o}\color{#ffaa00}{m}\color{#ffd400}{e}~\color{#bfff00}{t}\color{#80ff00}{o}~\color{#00ff00}{O}\color{#00ff40}{p}\color{#00ff80}{e}\color{#00ffbf}{n}\color{#00ffff}{S}\color{#00aaff}{t}\color{#0055ff}{u}\color{#0000ff}{d}\color{#2300ff}{y}\color{#4600ff}{!}\color{#6800ff}{!}\color{#8b00ff}{!}\\\small\cal Made~by~@TheSmartOne\) Hey there!!! Since you are new here, read this legendary tutorial for new OpenStudiers!! http://openstudy.com/study#/updates/543de42fe4b0b3c6e146b5e8

TheSmartOne (thesmartone):

Hmm. @satellite73 could become your best friend.. jk You know 73 73? Not a funny joke? lol ok then :)